OpenStudy (anonymous):

Find the exact circular function value tan-5pi/6 I need help

OpenStudy (anonymous):

a negative angle rotates clockwise...  since...\[\tan \left( \frac{ -5\pi }{ 6 } \right)=\frac{ \frac{ 1 }{ 2 } }{ \frac{ \sqrt{3} }{ 2 } }=\frac{ 1 }{ \sqrt{3} }\times \frac{ \sqrt{3} }{ \sqrt{3} }=\frac{ \sqrt{3} }{ 3 }\]